OpenStudy (anonymous):

Find the product of (x + 3)2. x2 + 6x + 9 x2 − 6x + 9 x2 − 9 x2 + 9

hartnn (hartnn):

\(\huge(a+b)^2=a^2+2ab+b^2 \\ \text{just put a=x and b=3, what u get ?}\)

OpenStudy (anonymous):

this one isnt the one i need help with. Its the one i just posted :/